You are conducting a test of the claim that the row variable and the column variable are dependent in the following contingency table.

X Y Z
A 38 23 30
B 34 33 31

Give all answers rounded to 3 places after the decimal point, if necessary.

(a) Enter the expected frequencies below:

X Y Z
A
B

(b) What is the chi-square test-statistic for this data?
      Test Statistic: χ2=χ2=

(c) What is the critical value for this test of independence when using a significance level of αα = 0.10?
      Critical Value: χ2=χ2=  

(d) What is the correct conclusion of this hypothesis test at the 0.10 significance level?

a)There is sufficient evidence to support the claim that the row and column variables are dependent.

b)There is not sufficient evidence to warrant rejection of the claim that the row and column variables are dependent.

c)There is not sufficient evidence to support the claim that the row and column variables are dependent.

d)There is sufficient evidence to warrant rejection of the claim that the row and column variables are dependent.
Answer #1
X Y Z Total
A 38 23 30 91
B 34 33 31 98
Total 72 56 61 189

a) E: Expected frequency.
E = ( Row total*Column total) / Grand total

X Y Z
A 34.667 26.963 29.370
B 37.333 29.037 31.630

b) Test statistic is

(O - E)2

O: Observed frequency
E: Expected frequency.

O E (O-E) (O-E)^2 (O-E)^2/E
38 34.667 3.333333 11.11111 0.320513
23 26.963 -3.96296 15.70508 0.582468
30 29.370 0.62963 0.396433 0.013498
34 37.333 -3.33333 11.11111 0.297619
33 29.037 3.962963 15.70508 0.540864
31 31.630 -0.62963 0.396433 0.012534
Total 1.77

c) Degrees of freedom = ( Number of rows - 1 ) * ( Number of column - 1) = ( 2 - 1) * (3 - 1) = 1 * 2 = 2

Level of significance = 0.10

Critical value = 4.602

d) Test statistic < critiical value we fail to reject null hypothesis.

Correct option is

c)There is not sufficient evidence to support the claim that the row and column variables are dependent.